# Incentivizing Reasoning for Advanced Instruction-Following of Large Language Models

### v4

@@ Abstract
Existing large language models (LLMs) face challenges of following complex instructions, especially when multiple constraints are present and organized in paralleling, chaining, and branching structures. One intuitive solution, namely chain-of-thought (CoT), is expected to universally improve capabilities of LLMs. However, we find that the vanilla CoT exerts a negative impact on performance due to its superficial reasoning pattern of simply paraphrasing the instructions. It fails to peel back the compositions of constraints for identifying their relationship across hierarchies of types and dimensions. To this end, we propose a systematic method to boost LLMs in dealing with complex instructions via incentivizing reasoning for test-time compute scaling. First, we stem from the decomposition of complex instructions under existing taxonomies and propose a reproducible data acquisition method. Second, we exploit reinforcement learning (RL) with verifiable rule-centric reward signals to cultivate reasoning specifically for instruction following. We address the shallow, non-essential nature of reasoning under complex instructions via sample-wise contrast for superior CoT enforcement. We also exploit behavior cloning of experts to facilitate steady distribution shift from fast-thinking LLMs to skillful reasoners. Extensive evaluations on seven comprehensive benchmarks confirm the validity of the proposed method, where a 1.5B LLM achieves 11.74% gains with performance comparable to a 8B LLM.
@@ 1 Introduction
Large language models (LLMs) exhibited remarkable performance on real-world tasks  Touvron et al. (2023); Grattafiori et al. (2024); Kojima et al. (2022); Zhao et al. (2023). Such generalizability is built upon the instruction-following capabilities of LLMs Ouyang et al. (2022); Mu et al. (2023); Zhou et al. (2023). To benchmark whether LLMs can produce the desired outputs under complex instructions, existing studies predominantly focus on modeling various types of constraints and rules where all requirements are expected to be satisfied simultaneously Zhang et al. (2024a); Jing et al. (2023); He et al. (2024a); Zhou et al. (2023). Recently, the compositions of constraints (And, Chain, Selection, and Nested) have been systemized Wen et al. (2024) to further enhance the complexity of instructions and demonstrate that LLMs still fail to meet expectations under intricate structures.
To improve LLMs in solving complex instructions, most prior methods leverage two kinds of techniques: 1) supervised fine-tuning (SFT) Xu et al. (2023); He et al. (2024b), and 2) template-guided inference Sun et al. (2024a); Chae et al. (2024). Despite their effectiveness, these methods are all task-specific and prone to overfitting of limited types of instructions. For the former, a large amount of curated instruction-response pairs are required to ensure diversity. For the latter, it is almost impossible to exhaustively enumerate the templates for problem decomposition, reflection, and refinement beforehand. Therefore, it calls upon a scalable solution that is both constructive and generalizable without meticulous manual efforts.
@@
In this paper, we propose to universally improve the instruction-following capabilities of LLMs under complex instructions. Inspired by the success of o1 Jaech et al. (2024), R1 Guo et al. (2025), and QwQ Qwen (2025), we resort to CoT for tackling complex instructions. We provide a practical guide for cultivation of effective reasoning tailored via reinforcement learning (RL). Specifically, our method addresses the following two main challenges: 1) the data scarcity of diverse complex instructions and 2) the secret recipe behind the formulation of effective CoT. For the former, although various data synthesis methods Xu et al. (2023); Wang et al. (2022a); Chung et al. (2024); Köpf et al. (2023); Tunstall et al. (2023); Chen et al. (2024); Luo et al. (2023); Sun et al. (2024a) have been proposed to pile up instructions, they do not take into consideration the taxonomy of constraint types and compositions. In contrast, we stem from the atomic rules, constraints, and their compositions to perform self-evolving across various tasks and domains. Besides, our instruction scaling is integrated with our RL-based reasoning stimulation. Both rule-based and model-based verification approaches, which serve as reward sources later, are paired with the instructions during generation. For the latter, existing approaches that seek to reproduce R1-like reasoning Hu et al. (2025); Pan et al. (2025); Luo et al. (2025); Zeng et al. (2025) are all restricted to solving maths problems with pure rule-based rewards, which are inapplicable for our settings. We consequently keep an eye on the differences between maths and complex instructions, and propose our RL recipe with rule-centric reward modeling, enforcement on superior CoT, and control on policy distribution drift. Our contributions are three-fold:
  • We propose a systematic method to boost LLMs under complex instructions via self-evolving instruction synthesis and reasoning-driven reinforcement learning.
@@
  • We validate our effectiveness on seven comprehensive benchmarks along with extensive generalization studies across model families, sizes, and cold/warm-start reasoners.
@@ 2.2 Complex Rule and Constraint Following
Numerous studies enhanced rule and constraint following capabilities via data engineering Qin et al. (2024b). The top concerns are around the synthesis of complicated constraints without costly human intervention. Specifically, WizardLM Xu et al. (2023) presents both in-breadth and in-depth evolving to generate complex instructions for tuning. Air Liu et al. (2025) stems from the simple instruction distilled from documents and then iteratively add one or more constraints to increase complexity. Back-translation Qi et al. (2024) is another popular technique to introduce additional constraints. Self-play Dong et al. (2024b) targets at generating constraints that can be verified by codes. SPAR Cheng et al. (2024) combines both self-play and tree-search for iterative refinement. The detailed taxonomy of rules and constraints Zhang et al. (2024a); Wang et al. (2024a); Sun et al. (2024a), together with their structures and relationship Hayati et al. (2024); Wen et al. (2024), allows precise synthesis control over instruction types. With respect to the post-training technique, most methods adopt the SFT Sun et al. (2024b) and RL Huang et al. (2025a); Wang et al. (2024b); Zhang et al. (2024b); Wu et al. (2024b); Ren et al. (2025) for preference alignment. Moreover, training-free workflows have also been proposed to decompose the instructions  Ferraz et al. (2024) and exploit verifiable feedback Wang et al. (2025) for refinement until all the check-boxes are ticked Cook et al. (2024).
@@ 3.1 Problem Definition
Our goal is to incentivize reasoning of a LLM to advance its capabilities of handling and solving complex instructions (see Fig. 2). Through systematic pipeline investigation and extensive experimental analyses, we offer key insights and practical guidelines to arm LLMs with developed reasoning.
@@
3.2 Self-Evolving Instructions with Complex Rules and Constraints
To address the scarcity of complex instructions, we propose to scale up instructions with various rules, constraints and their verification criteria via self-evolving.
@@ 3.2 Self-Evolving Instructions with Complex Rules and Constraints
Self-Instruct with Rules and Constraints
We adapt self-instruct Wang et al. (2022a) under different fine-grained rules and constraints Zhang et al. (2024a). In view of the verification techniques, both code-execution Qiao et al. (2023) and LLM-as-a-Judge Zheng et al. (2023) are involved to provide evaluation feedback. For the former, we prepare a collection of constraint templates and their executable codes. Then, we randomly instantiate a combination of rules and constraints from the pool. To ensure their mutual compatibility, a pre-defined validity check is enforced to eliminate conflicts (e.g., The first paragraph must start with... and Wrap your entire reply with...). For the latter, we construct pairs of sub-instructions and scoring questions that stress on the style and semantic constraints. Such sub-instructions correspond to constraints with LLM-based evaluation (e.g., Is the answer written in the tone of Confucious?), which are complementary to those relying on code-based evaluation. An off-the-shelf LLM is utilized to perform few-shot in-breadth evolution. With respect to the composition of these atomic sub-instructions, we refer to He et al. (2024b) for definitions of And, Chain, and Selection. These sub-instructions are assembled for the integrated instructions. Details can be found in the Sec. A.3.2.
@@
We use LLMs to generate responses and filter out low-quality query-response pairs that fail to pass the associated verification tests. Additionally, we observe that the self-evolved instructions still contain unreasonable constraints or nonsensical queries (e.g., Give me a very short, concise, and clear response...The response should have 4 sections.). Under such circumstance, we summarize seven typical issues and curate judgment prompts for LLMs to double-check the instructions and keep the valid ones (see Sec. A.3.3).
@@ 4.1 Experimental Setup
Details on statistics about our self-evolving dataset can be found in Sec. A.3. Additionally, we incorporated the DeepScaleR Luo et al. (2025) maths dataset to facilitate reasoning development (see Sec. A.4). The evaluation metrics for each benchmark are reported in Sec. A.2.2.
@@
We use OpenRLHF Hu et al. (2024) for both cold-start (Qwen2.5-1.5B/7B Yang et al. (2024c), LLaMA3.1-8B Grattafiori et al. (2024), and Ministral-8B Jiang et al. (2023b)) and warm-start (DeepSeek-Qwen1.5B/7B Guo et al. (2025) and DeepScaleR-1.5B Luo et al. (2025)) experiments. Detailed settings can be found in Sec. A.5. To ensure fairness and comparability, we employ Qwen2.5-72B-Instruct as the judge for all LLM-based evaluation.
@@
LLaMA3.1-8B-Instruct I/O 77.63 75.20 56.99 69.11 46.92 53.52 71.52 67.01
LLaMA3.1-8B-Instruct CoT 60.44 65.50 47.66 56.54 32.34 37.36 58.48 54.53(-12.48%)
LLaMA3.1-8B-Instruct SDC 80.22 71.00 58.33 68.73 38.36 48.92 72.89 65.24(-1.77%)
LLaMA3.1-8B-Instruct SFT 77.26 75.80 54.00 65.24 40.16 59.56 65.30 64.92(-2.09%)
LLaMA3.1-8B-Instruct Ours 13.49 4.6 1.33 2.71 7.14 1.08 0.51 4.06(-62.95%)
@@ 4.2 Main Results
Our method effectively boosts most of the existing LLMs in handling complex instructions (see Table 1), demonstrating the generalization of the cultivated deep reasoning. In contrast, the CoT prompting causes a drastic performance decline to all models, which further confirms the detrimental effect of shallow thinking. Instead of adopting an one-off generation, SDC decouples the reasoning and answering via two-step inference. Due to the intrinsic superficial nature, SDC still fails to improve the reasoning quality. The SFT technique directly performs knowledge distillation where small LLMs mimic the reasoning patterns of strong slow-thinker. It guarantees that the depth and breadth of thinking is under immediate supervision. However, one drawback of SFT is that the model’s generalization ability tends to deteriorate for samples that fall outside the domains encountered during training. Comparatively, our RL paradigm teaches LLMs how to think, driving the self-development of varied reasoning rather than simple memorization.
In line with Fig. 3, small models (1.5B) achieve much more gains than larger ones, showcasing the potentials of small LLMs via test-time scaling. The DeepSeek-distilled LLMs possess a good starting point for reasoning organization from their warm-start imitation across a broad range of tasks and topics (Fig. 36). With respect to model family, we unfortunately find that the capacity of Ministral and LLaMA is inferior to that of Qwen. The Ministral-8B exhibits limited advantages over its vanilla counterpart while the LLaMA3.1-8B experienced a model collapse during training. As shown in Fig. 35(a) and (d), a rapid shrinkage of response and a frequent surge of KL penalty imply a great deviation of LLaMA from its initial state. The reason behind might be ascribed to the pre-trained knowledge of base models Gandhi et al. (2025). LLaMA tends to generate endless thinking without conforming to the required format. It then struggles to output semantically consistent responses and keeps extending its meaningless thinking until collapse. Detailed results can be found in Sec. A.7.1. Discussions on generalization over multi-purpose benchmarks can be found in Sec. A.7.2.
@@
We implemented SOTAs on the ComplexBench (see Table 2): oracle decomposition Wen et al. (2024) (ground-truth sub-instructions), self-consistency Wang et al. (2022b) (majority voting@10), Conifer Sun et al. (2024a) (SFT), FollowComplex He et al. (2024b) (SFT), and Self-Refine Madaan et al. (2023). Our method demonstrate its superiority on the most complicated Chain, Selection categories, suggesting that the reasoning indeed assists analysis of LLMs to carry out the truly relevant, valid request with constraints.
@@
Table 3: Ablation study on the Qwen2.5-7B-Instruct with CoT reasoning. Best/2nd best are marked bold/underlined. Maths and Complex refer to the DeepScaleR and our self-evolved dataset, respectively. SupCoT and BC denote the superior CoT enforcement and behavior cloning, respectively.

### v5

@@ Abstract
Existing large language models (LLMs) face challenges of following complex instructions, especially when multiple constraints are present and organized in paralleling, chaining, and branching structures. One intuitive solution, namely chain-of-thought (CoT), is expected to universally improve capabilities of LLMs. However, we find that the vanilla CoT exerts a negative impact on performance due to its superficial reasoning pattern of simply paraphrasing the instructions. It fails to peel back the compositions of constraints for identifying their relationship across hierarchies of types and dimensions. To this end, we propose RAIF, a systematic method to boost LLMs in dealing with complex instructions via incentivizing reasoning for test-time compute scaling. First, we stem from the decomposition of complex instructions under existing taxonomies and propose a reproducible data acquisition method. Second, we exploit reinforcement learning (RL) with verifiable rule-centric reward signals to cultivate reasoning specifically for instruction following. We address the shallow, non-essential nature of reasoning under complex instructions via sample-wise contrast for superior CoT enforcement. We also exploit behavior cloning of experts to facilitate steady distribution shift from fast-thinking LLMs to skillful reasoners. Extensive evaluations on seven comprehensive benchmarks confirm the validity of the proposed method, where a 1.5B LLM achieves 11.74% gains with performance comparable to a 8B LLM. Evaluation on OOD constraints also confirms the generalizability of our RAIF.
@@ 1 Introduction
Large language models (LLMs) exhibited remarkable performance on real-world tasks  Touvron et al. (2023); Grattafiori et al. (2024); Kojima et al. (2022); Zhao et al. (2023). Such generalizability is built upon the instruction-following capabilities of LLMs Ouyang et al. (2022); Mu et al. (2023); Zhou et al. (2023). To benchmark whether LLMs can produce the desired outputs under complex instructions, existing studies predominantly focus on modeling various types of constraints and rules where all requirements are expected to be satisfied simultaneously Zhang et al. (2024a); Jing et al. (2023); He et al. (2024a); Zhou et al. (2023). Recently, the compositions of constraints (And, Chain, Selection, and Nested) have been systemized Wen et al. (2024) to further enhance the complexity of instructions and demonstrate that LLMs still fail to meet expectations under intricate structures. These complex instructions, which are often composed of multiple sub-instructions, enforce various constraints on the expected responses. Existing LLMs either ignore certain constraints in And structures or misinterpret the instruction to respond to the wrong sub-instructions in Selection structures.
To improve LLMs in solving complex instructions, most prior methods leverage two kinds of techniques: 1) supervised fine-tuning (SFT) Xu et al. (2023); He et al. (2024b), and 2) template-guided inference Sun et al. (2024a); Chae et al. (2024). Despite their effectiveness, these methods are all task-specific. They are prone to overfitting constraints in the training set and fail to generalize to the unseen ones Lambert et al. (2024). For the former, a large amount of curated instruction-response pairs are required to ensure diversity. For the latter, it is almost impossible to exhaustively enumerate the templates for problem decomposition, reflection, and refinement beforehand. Therefore, it calls upon a scalable solution that is both constructive and generalizable without meticulous manual efforts.
@@
In this paper, we propose RAIF to leverage Reasoning for Advancing the Instruction-Following capabilities of LLMs under the context of complex instructions. Inspired by the success of o1 Jaech et al. (2024), R1 Guo et al. (2025), and QwQ Qwen (2025), we resort to CoT for tackling complex instructions. We provide a practical guide for cultivation of effective reasoning tailored via reinforcement learning (RL). Specifically, our method addresses the following two main challenges: 1) the data scarcity of diverse complex instructions and 2) the secret recipe behind the formulation of effective CoT. For the former, although various data synthesis methods Xu et al. (2023); Wang et al. (2022a); Chung et al. (2024); Köpf et al. (2023); Tunstall et al. (2023); Chen et al. (2024); Luo et al. (2023); Sun et al. (2024a) have been proposed to pile up instructions, they do not take into consideration the taxonomy of constraint types and compositions. In contrast, we stem from the atomic rules, constraints, and their compositions to perform LLM-based evolving across various tasks and domains. Besides, our instruction scaling is integrated with our RL-based reasoning stimulation. Both rule-based and model-based verification approaches, which serve as reward sources later, are paired with the instructions during generation. For the latter, existing approaches that seek to reproduce R1-like reasoning Hu et al. (2025); Pan et al. (2025); Luo et al. (2025); Zeng et al. (2025a) are all restricted to solving maths problems with pure rule-based rewards, which are inapplicable for our settings. We consequently keep an eye on the differences between maths and complex instructions, and propose our RL recipe with rule-centric reward modeling, enforcement on superior CoT, and control on policy distribution drift. Our contributions are three-fold:
  • We propose a systematic method to boost LLMs under complex instructions via LLM-based evolving for instruction synthesis and reasoning-driven reinforcement learning.
@@
  • We validate our effectiveness on seven comprehensive benchmarks along with extensive generalization studies across model families, sizes, cold/warm-start reasoners, and OOD constraints.
@@ 2.2 Complex Rule and Constraint Following
Numerous studies enhanced rule and constraint following capabilities via data engineering Qin et al. (2024b). The top concerns are around the synthesis of complicated constraints without costly human intervention. Specifically, WizardLM Xu et al. (2023) presents both in-breadth and in-depth evolving to generate complex instructions for tuning. Air Liu et al. (2025) stems from the simple instruction distilled from documents and then iteratively add one or more constraints to increase complexity. Back-translation Qi et al. (2024) is another popular technique to introduce additional constraints. Self-play Dong et al. (2024b) targets at generating constraints that can be verified by codes. SPAR Cheng et al. (2024) combines both self-play and tree-search for iterative refinement. The detailed taxonomy of rules and constraints Zhang et al. (2024a); Wang et al. (2024a); Sun et al. (2024a), together with their structures and relationship Hayati et al. (2024); Wen et al. (2024); Zeng et al. (2025b), allows precise synthesis control over instruction types. With respect to the post-training technique, most methods adopt the SFT Sun et al. (2024b) and RL Huang et al. (2025a); Wang et al. (2024b); Zhang et al. (2024b); Wu et al. (2024b); Ren et al. (2025) for preference alignment. Moreover, training-free workflows have also been proposed to decompose the instructions  Ferraz et al. (2024) and exploit verifiable feedback Wang et al. (2025) for refinement until all the check-boxes are ticked Cook et al. (2024). The most recent recipe of post-training by Tülu 3 Lambert et al. (2024) augments the existing instructions with IFEval constraints and optimizes the LLM with RL for better alignment. The differences between Tülu 3 and ours are fourfold: 1) Problem definition. We aim at solving a broad range of complex instructions with various constraints and their compositions via reasoning. Tülu 3 focuses only on responding to the IFEval-style instructions without the preceding reasoning, and therefore exhibits limited generalization on out-of-domain (OOD) constraints. 2) Constraint type and structure. We follow CFBench [8] and ComplexBench [11] respectively for the constraint categories and composition types. Tülu 3 is limited in code-verifiable constraints with ”and” structures. 3) Reward modeling. Both code-execution and LLM-as-a-Judge are involved for verification. This greatly expands the tasks that can be handled by our method, where the semantic constraints with LLM-as-a-Judge verifications are complementary to the lexical and word constraints. 4) RL algorithm. Tulu 3 directly optimizes the response via the vanilla PPO algorithm. However, we are aimed at incentivizing reasoning for solving complex instructions. Since LLMs are prone to shallow reasoning and tend to muddle with a few sentences that ignore the constraints and their structures, we design the filtering mechanism that only selects responses with deep, true reasoning. Our proposed superior CoT enforcement acts as a bridge between reasoning and instruction-following quality.
@@ 3.1 Problem Definition
Our goal is to incentivize reasoning of a LLM to advance its capabilities of handling complex instructions (see Fig. 2). Through systematic pipeline investigation and extensive experimental analyses, we offer key insights and practical guidelines to arm LLMs with deep reasoning. Our research scope focuses on instructions that consist of one or more atomic sub-instructions with compositional structures (And, Selection, Chain, and Nested). Multiple constraints are enforced so that LLMs have to carefully comprehend the instructions and reason on which sub-instructions to perform execution and how to obey all the rules.
@@
3.2 LLM-based Instruction Evolving with Complex Rules and Constraints
To address the scarcity of complex instructions, we propose to scale up instructions with various rules, constraints and their verification criteria via LLM-based evolving.
@@ 3.2 LLM-based Instruction Evolving with Complex Rules and Constraints
Instruction Evolving with Rules and Constraints
We adapt self-instruct Wang et al. (2022a) for instruction evolving under different fine-grained rules and constraints Zhang et al. (2024a). In view of the verification techniques, both code-execution Qiao et al. (2023) and LLM-as-a-Judge Zheng et al. (2023) are involved to provide evaluation feedback. For the former, we prepare a collection of constraint templates and their executable codes. Then, we randomly instantiate a combination of rules and constraints from the pool. To ensure their mutual compatibility, a pre-defined validity check is enforced to eliminate conflicts (e.g., The first paragraph must start with... and Wrap your entire reply with...). For the latter, we construct pairs of sub-instructions and scoring questions that stress on the style and semantic constraints. Such sub-instructions correspond to constraints with LLM-based evaluation (e.g., Is the answer written in the tone of Confucious?), which are complementary to those relying on code-based evaluation. An off-the-shelf LLM is utilized to perform few-shot in-breadth evolution. With respect to the composition of these atomic sub-instructions, we refer to He et al. (2024b) for definitions of And, Chain, and Selection. These sub-instructions are assembled for the integrated instructions. Details can be found in the Sec. A.3.2.
@@
We use LLMs to generate responses and filter out low-quality query-response pairs that fail to pass the associated verification tests. Additionally, we observe that the LLM-evolved instructions still contain unreasonable constraints or nonsensical queries (e.g., Give me a very short, concise, and clear response...The response should have 4 sections.). Under such circumstance, we summarize seven typical issues and curate judgment prompts for LLMs to double-check the instructions and keep the valid ones (see Sec. A.3.3).
@@ 3.3 Reinforcement Learning for Reasoning Under Complex Instructions
The replay buffer filtering guarantees that every retained CoT is causally linked to better instruction-following performance rather than just longer text. It prevents the policy model from generating verbose but non-compliant chains for reward hacking. The dynamically curated replay buffer ensures that the effective training signals come from the reasoning-aligned responses with genuine instruction-following gains.
@@ 4.1 Experimental Setup
Details on statistics about our LLM-evolved dataset can be found in Sec. A.3. Additionally, we incorporated the DeepScaleR Luo et al. (2025) maths dataset to facilitate reasoning development (see Sec. A.4). The evaluation metrics for each benchmark are reported in Sec. A.2.2.
@@
We use OpenRLHF Hu et al. (2024) for both cold-start (Qwen2.5-1.5B/7B Yang et al. (2024c), and Ministral-8B Jiang et al. (2023b)) and warm-start (DeepSeek-Qwen1.5B/7B Guo et al. (2025) and DeepScaleR-1.5B Luo et al. (2025)) experiments. Detailed settings can be found in Sec. A.5. To ensure fairness and comparability, we employ Qwen2.5-72B-Instruct as the judge for all LLM-based evaluation.
@@ 4.2 Main Results
Our method effectively boosts most of the existing LLMs in handling complex instructions (see Table 1), demonstrating the generalization of the cultivated deep reasoning. In contrast, the CoT prompting causes a drastic performance decline to all models, which further confirms the detrimental effect of shallow thinking. Instead of adopting an one-off generation, SDC decouples the reasoning and answering via two-step inference. Due to the intrinsic superficial nature, SDC still fails to improve the reasoning quality. The SFT technique directly performs knowledge distillation where small LLMs mimic the reasoning patterns of strong slow-thinker. It guarantees that the depth and breadth of thinking is under immediate supervision. However, one drawback of SFT is that the model’s generalization ability tends to deteriorate for samples that fall outside the domains encountered during training. Comparatively, our RL paradigm teaches LLMs how to think, driving the development of varied reasoning rather than simple memorization. We believe that SFT might be a good starting point, but it does not promise generalization.
In line with Fig. 3, small models (1.5B) achieve much more gains than larger ones, showcasing the potentials of small LLMs via test-time scaling. With respect to the 7B-scale models, although their performance gains are smaller than 1.5B models, the consistent improvement across cold-start and warm start models are still non-trivial on seven benchmarks. For larger models, the current dataset size ( 13K) might not be enough. With an increase of dataset size, 7B-scale models are expected to achieve superior performance according to the scaling law Team et al. (2025); Guha et al. (2025); Lambert et al. (2024). In the future, we will construct mixed RLVR dataset with size over 100K for scaling up.
The DeepSeek-distilled LLMs possess a good starting point for reasoning organization from their warm-start imitation across a broad range of tasks and topics (Fig. 35). With respect to model family, we unfortunately find that the capacity of Ministral is inferior to that of Qwen. The Ministral-8B exhibits limited advantages over its vanilla counterpart. Detailed results can be found in Sec. A.7.1. Discussions on generalization over multi-purpose benchmarks can be found in Sec. A.7.2.
@@
Table 3: Training and testing-time compute on ComplexBench (Qwen2.5-7B-Instruct). OD, SC, CNFR, FC, and SR stand for the Oracle Decomposition Wen et al. (2024), Self-Consistency Wang et al. (2022b), Conifer Sun et al. (2024a), FollowComplex He et al. (2024b), and Self-Refine Madaan et al. (2023).
Method Training Compute Test-Time Compute $\delta$
# SFT
samples
# SFT
epochs
RLHF
algo.
# RLHF
samples
# RLHF
epochs
Avg. # steps
or sampling
Avg. #
reasoning tokens
Avg. #
answer tokens
Avg. gains
over I/O
I/O 1 332 74.47
CoT 1 80 419 -2.46
SDC 1 79 372 +2.06
SFT 13K 10 1 361 -0.23
OD [11] 1 407 +1.79
SC [98] 10 3302 -0.70
SR [99] 1.9 877 -0.46
CNFR [14] 66K 4 DPO 63K 1 1 308 -10.96
FC [13] 12K 2 DPO 12K 2 1 261 -2.50
Ours GRPO 26K 3 1 299 349 +2.93
@@
We implemented SOTAs on the ComplexBench (see Table 2): Oracle Decomposition Wen et al. (2024) (ground-truth decomposition of sub-instructions), Self-Consistency Wang et al. (2022b) (majority voting@10), Conifer Sun et al. (2024a), FollowComplex He et al. (2024b), and Self-Refine Madaan et al. (2023). We also report the training and testing-time compute in Table 3. Our method demonstrate its superiority on the most complicated Chain, Selection categories, suggesting that the reasoning indeed assists analysis of LLMs to carry out the truly relevant, valid request with constraints.
Due to the nature of shallow reasoning of instructed LLMs, most training-free methods that use manually designed workflows (e.g., Self-Consistency and Self-Refine) do not benefit from the increased test-time compute. Despite the maximum six interaction turns, Self-Refine tends to terminate at the second turn as LLMs are over-confident about their own initial responses and fail to detect unsatisfied constraints for correction. The SDC method, by decoupling reasoning from answer generation, improves performance over the vanilla CoT with a similar test-time compute budget. The gains of oracle decomposition confirm that the removal of redundant sub-instructions and constraints is beneficial. However, such decomposition requires human annotation and is not practical. On the other hand, Conifer Sun et al. (2024a) and FollowComplex He et al. (2024b) employ both SFT and direct preference optimization (DPO) where the meticulous selection of positive and negative pairs is indispensable. Our RAIF shares a comparable training compute with Conifer and FollowComplex, but achieves a superior performance due to the generalization of deep reasoning on diverse constraints and structures. In consideration of our gains over I/O, the slight increase in test-time compute is acceptable (i.e., 299 tokens).
@@
Table 4: Ablation study on the Qwen2.5-7B-Instruct with CoT reasoning. Best/2nd best are marked bold/underlined. Maths and Complex refer to the DeepScaleR and our LLM-evolved dataset, respectively. SupCoT and BC denote the superior CoT enforcement and behavior cloning, respectively.